The ratio of the number of sides of a square to the number of edges of a cube is
(A) 1 : 2 (B) 3 : 2 (C) 4 : 1 (D) 1 : 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (D) is correct here.

We know that, number of sides in a square = 4 and number of edges in a cube = 12 So, ratio of sides to edges = 4/12
Divide both numerator and denominator by 4.

= 1/3
Option (D) is correct here.

We know that, number of sides in a square = 4 and number of edges in a cube = 12
So, ratio of sides to edges = 4/12
Divide both numerator and denominator by 4.

= 1/3
Therefore, ratio of sides to edges = 1 : 3

Thanks